OpenStudy (lilcathyx):

I just need someone to check :/ Determine if triangle RST with coordinates R (2, 3), S (4, 4), and T (5, 0) is a right triangle. Use evidence to support your claim. If it is not a right triangle, what changes can be made to make it a right triangle? Be specific

OpenStudy (lilcathyx):

This is my answer R(2, 3) & T(5, 0) m= y2-y1 / x2-x1 m= 0-3 / 5-2 m= -3 /3 = -1 S(4, 4) & T(5, 0) m= y2-y1 / x2-x1 m= 0-4 / 5-4 m= -4 / 1 This is not a right triangle

OpenStudy (lilcathyx):

@IMStuck @ParthKohli @ganeshie8 @Ankh @iambatman

OpenStudy (lilcathyx):

What do you think ? o:

OpenStudy (anonymous):

You found the slopes of ST and RT, how did they help you prove it's not a right triangle?

OpenStudy (lilcathyx):

they are not opposite reciprocals

OpenStudy (anonymous):

True, but what about RS?

OpenStudy (anonymous):

What if it's opposite to SR and RT?

OpenStudy (anonymous):

When you're given coordinates it's recommended you plot them to get a visual perspective on the shape.

OpenStudy (anonymous):



OpenStudy (lilcathyx):

RS = 1/2

OpenStudy (anonymous):

When you do a quick sketch you can see that there's no way 90 degrees would fit between ST and RT. The other two angles are questionable though, so you need the slope of RS as well.

OpenStudy (anonymous):

Once you find all 3 slopes, only then you can say that none of them are opposite reciprocals.

OpenStudy (lilcathyx):

Mhmmm. so all i have to add to my answer is that. if i do a sketch of the coordinates, there is no 90 degree angle ? or i have to add the slope of RS and

OpenStudy (lilcathyx):

ohhhhh

OpenStudy (lilcathyx):

so now i would put the slope of RS and write This triangle is not a right triangle because it does not have opposite reciprocals ?

OpenStudy (anonymous):

Yes, because out of the three slopes there is no pair that are opposite reciprocals.

OpenStudy (lilcathyx):

Oooohh okkay thank you ^.^ can u help me in 1 more ?

OpenStudy (anonymous):

The second part?

OpenStudy (lilcathyx):

no, if you could help me in another question ? i promise its just 1

OpenStudy (anonymous):

Sure.